# Topological characterization of magnon-polaron bands and thermal Hall conductivity in a frustrated kagome antiferromagnet

### v1

@@ III.3 Magnon-polaron thermal Hall conductivity
Additionally, in Fig. 6(c), we have also shown the direct dependence of thermal Hall conductivity ($\kappa_{xy}$) on temperature at three distinct values of $g^{l}$ (a $g^{l}=0$ is included for comparison). In the absence of spin-phonon coupling, the Hall conductivity arises purely from magnonic states. With no spin-phonon coupling at low temperatures, $\kappa_{xy}$ acquires a positive contribution and reaches a local maximum, beyond which it decreases and eventually becomes negative. It may be noted that, at a finite temperature, $\kappa_{xy}$ vanishes due to the mutual cancellation of contributions from all the bands. Such behavior of the thermal Hall conductivity is well-established in the literature [102, 20]. Additionally, at very high temperatures, $\kappa_{xy}$ tends to saturate to a finite value.
Figure 7: (a) The Chern number corresponding to the lower and the middle bands is plotted. Within the dotted region, there occurs closure of gaps between the respective bands. (b) The bulk band structure illustrates the gap-closing transition. The Berry curvature plots corresponding to the lower band, (c) before and (d) after the gap closing transition, are displayed.
Figure 8: The behavior of the thermal Hall conductivity ($\kappa_{xy}$) in the presence of non-local spin-phonon interaction is shown. (a) The thermal Hall conductivity is plotted as a function of external magnetic field for different values of the coupling strength $g^{nl}$ at $\widetilde{T}=0.4$. The transition lines $\bar{s}^{1},~{}\bar{s}^{2}$ denote $g^{nl}=0.4$, while $\bar{s}^{3},~{}\bar{s}^{4}$ refer to $g^{l}/J_{1}=0.8$, and the $s^{0}$ transition line corresponds to a topological phase transition for all the cases. (b) Thermal Hall conductivity is also presented as a function of $g^{nl}$ at two different temperatures, keeping $B_{0}=0.4B_{s}$, where $\bar{v}^{1}$ and $\bar{v}^{2}$ correspond to $\widetilde{T}=0.4$ and $0.8$, respectively. Additionally, (c) the thermal Hall conductivity is shown as a function of temperature for distinct values of $g^{nl}$. Here, $\widetilde{\kappa}_{xy}=\kappa_{xy}/J_{1}S$ and $\widetilde{T}=k_{B}T/J_{1}S$. Throughout this analysis, the other parameters are fixed at $J_{2}=0.03J_{1}$ and $D=0.045J_{1}$.
In the presence of local spin-phonon coupling ($g^{l}\neq 0$), the Hall conductivity corresponding to larger values of $g^{l}$ shows a substantial enhancement at low temperatures, whereas at higher temperatures, the growth becomes comparatively less pronounced as seen for $g^{l}=0.8J_{1}$ compared to $g^{l}=0.4J_{1}$ in Fig. 6(c). While in the absence of spin-phonon coupling, the magnon energy bands are independent of temperature, in its presence, the magnon-polaron states are no longer independent of temperature. As temperature increases, we observe positive energy shifts corresponding to the two lower bands (see Apendix C), which affects the $c_{2}(\rho_{n,\mathbf{k}})$ function in such a way that the overall contribution from $c_{2}(\rho_{n,\mathbf{k}})$ function along with temperature responsible for the reduced slope in the behavior of the thermal Hall conductivity beyond a certain range as it increases. Notably, at higher values of the coupling strength, this behavior is more prominent. This type of phenomenon reflects the suppression of thermal Hall conductivity due to phonons at high temperatures.
Further, for $g^{l}=0.4J_{1}$, we find a topological phase transition from $C=(1,-2,1)$ to $C=(-1,0,1)$ occurring at the transition line $T_{c}$. For clarity, the corresponding changes in the Chern numbers of the participating bands are presented in Fig. 7(a) as a function of temperature ($\widetilde{T}$). Moreover, we observe that the effects of temperature are sufficiently weak such that the gap-closing transition at the $\mathbf{K}$ point (see Fig. 7(b)) extends over a broad temperature range. This region is highlighted by the dotted area in Fig. 7(a), signifying a weakly semimetalic phase of the bosonic quasiparticles. Owing to the gradual nature of the transition and the fact that it occurs at the $\mathbf{K}$ point, where the contribution from $c_{2}(\rho_{n,\mathbf{k}})$ is relatively small, the transition is not prominently noticeable from the Hall conductivity plot in Fig. 6(c). Furthermore, in Fig.7(c), for the lower band, we show the corresponding Berry curvature prior to the gap-closing transition, where positive values of the Berry curvature are concentrated at the high-symmetry $\mathbf{K}$ and $\mathbf{K^{\prime}}$ points, and they become negative beyond the gap-closing transition in Fig.7(d). This indicates a topological phase transition denoted by $C=1\to C=-1$.
Similar to the local interaction case, non-local spin-phonon interactions can also lead to some interesting topological phase transitions. The variation of the $\kappa_{xy}$ can be studied as a function of the external magnetic field ($B_{0}$) for two representative values of $g^{nl}$, while keeping the other parameters fixed at $J_{2}=0.03J_{1}$, $D=0.045J_{1}$, $\widetilde{T}=0.4$. For a direct comparison, we also include the results for the bare magnon case ($g^{l(nl)}=0$), as shown in Fig. 8(a).
At finite $g^{nl}$, namely, $g^{nl}=0.4J_{1}$, a gradual increase in $B_{0}$ leads to a phase transition from $C=(3,-1,-2)$ to $C=(3,-4,1)$ across the $\bar{s}^{1}$ transition line at $B_{0}=0.12B_{s}$, where the Hall conductivity reaches its peak value. A further increase in $B_{0}$ induces another phase transition at $\bar{s}^{2}$ at $B_{0}=0.25B_{s}$, where the Hall conductivity monotonically decreases. A similar sequence of phase transitions is observed for $g^{nl}=0.8J_{1}$, as shown in Fig. 8(a), where the transition lines $\bar{s}^{3}$ and $\bar{s}^{4}$ are obtained for $B_{0}=0.18B_{s}$ and $B_{0}=0.3B_{s}$, respectively. Moreover, in contrast to the earlier scenario involving local spin-phonon interactions, where the Hall conductivity for different values of $g^{l}$ changes sign (from positive to negative) at distinct value of $B_{0}$, the case corresponding to the non-local interaction exhibits nearly a constant $B_{0}$ at which the Hall conductivity vanishes for $g^{nl}=0.4J_{1}$ and $0.8J_{1}$ (see the clustering of the points on the horizontal dashed line). Additionally, near the $s^{0}$ transition line, a phase transition from $C=(1,-2,1)$ to $C=(-1,0,1)$ occurs for both $g^{nl}=0.4J_{1}$ and $0.8J_{1}$, as shown in Fig. 8(a).
Further, in Fig. 4(a) we have seen that at $\widetilde{T}=0$, with the same values of $D,~{}J_{2},~{}B_{0}$ mentioned earlier, one phase transition occurs from $C=(3,-4,1)$ to $C=(1,-2,1)$ as a function of $g^{nl}$ near $g^{nl}=0.43J_{1}$. In Fig. 8(b), we have shown the corresponding nature of the thermal Hall conductivity as a function of $g^{nl}$ at two different temperatures, where we observe a slight non-monotonic behavior of the Hall conductivity associated with that phase transition across the transition line $\bar{v}^{1}$ (at $g^{nl}=0.35J_{1}$) and $\bar{v}^{2}$ (at $g^{nl}=0.31J_{1}$) corresponding to the temperatures $\widetilde{T}=0.4$ and $0.5$, respectively. Moreover, it reflects how the transition points shift with varying temperature. Similar to the local case, the Hall conductivity increases with the coupling strength $g^{nl}$ except in the vicinity of the phase transition, as shown in Fig. 8(b).
Additionally, we have also shown the behavior of Hall conductivity as a function of temperature. Here we keep $B_{0}=0.4B_{s}$ same as in Fig. 6(c), and if we see Fig. 8(a), we can observe near $B_{0}=0.4B_{s}$, the behavior of the Hall conductivity remains nearly same for all the cases corresponding to different $g^{nl}$ which is manifested Fig. 8(c). Moreover, there are a couple of crossings between the Hall conductivity curves corresponding to $g^{nl}=0.4$ (in pink) and $0.8$ (in blue) as seen in Fig. 8(c). The first crossing at low temperature (near $\widetilde{T}\sim 0.3$) occurs due to a phase transition from $C=(3,-4,1)$ to $C=(1,-2,1)$, corresponding to the case with $g^{nl}=0.4J_{1}$, near the transition line $\overline{T}_{c}$ at $\widetilde{T}=0.29$. A second crossing exists at higher values of $\widetilde{T}$, which however is due to the change in the slope of $\kappa_{xy}$ corresponding to the case with $g^{nl}=0.8J_{1}$, which was earlier denoted as the suppression of the thermal Hall conductivity at high temperature induced by phonons.
@@
Figure 9: (a) The Chern numbers corresponding to the lower and the middle band are shown. Within the dotted region, there occurs closure of gaps between the respective bands. (b) The bulk band structure illustrates the gap-closing transition. The Berry curvature plots corresponding to the lower band, (c) before and (d) after the gap closing transition, are shown.
@@
For further clarity, we have presented the corresponding Chern number plot as a function of temperature in Fig. 9 (a), where the dotted region in light green background corresponds to the spectral gap-closing regime. Further, in Fig. 9(b), we observe a gap-closing transition near the $\mathbf{\Gamma}$ point, where the contribution from $c_{2}(\rho_{n,\mathbf{k}})$ is significant at low temperatures. As a result, the transition appears more prominently in the Hall conductivity plot compared to the case with local spin-phonon interaction. Additionally, in Fig. 9(c), at $\widetilde{T}=0.24$, prior to the gap-closing transition, we show the Berry curvature of the lower band. Positive values of the Berry curvature are observed near both the high-symmetry points ($\mathbf{\Gamma}$ and $\mathbf{K}$), corresponding to a Chern number of $C=3$. In contrast, in Fig. 9(d), after closing the spectral gap, the Berry curvature changes sign (negative) at the gap-closing point $\mathbf{\Gamma}$, while remaining positive at the $\mathbf{K}$ point, indicating a Chern number of $C=1$.
@@ IV CONCLUSION
We investigate the effects of couplings between spins and optical phonons in a topologically protected magnetic system, namely, in a frustrated Kagome antiferromagnet. Specifically, we consider two different types of spin-phonon couplings, one in which the interaction is purely local, while the other case accounts for non-local interactions. To derive the hybridized magnon-polaron bands, we perform detailed and systematic calculations based on a canonical spin-Peierls transformation formulated within the magnon framework. Interestingly, we observe several topological phase transitions solely induced by the spin-phonon coupling. The signatures for these topological transitions are explored via the magnon-polaron band structures, their Chern numbers, and the corresponding phase plots. The thermal Hall conductivities relevant to these scenarios are obtained, thereby providing further support for these topological transitions.
Let us enumerate the highlights of the results obtained by us. In the presence of both local and non-local spin-phonon couplings, we observe multiple topological phase transitions as a function of the DMI and the spin-phonon coupling strength. To establish the bulk boundary correspondence, the associated edge modes for distinct topological phases are computed on ribbon geometries, where we observe that a constant value of the coupling strength can lead to distinct topological phases depending on whether the coupling is local or non-local. Additionally, at finite temperatures, the thermal Hall conductivity plots as a function of the external magnetic field, spin-phonon coupling strength, and temperature effectively capture all the topological phases and phase transitions occurring therein. Furthermore, in this study, we observe that due to the inclusion of phonon contributions at finite-temperature, the spectral properties of the energy bands depend on temperature, inducing topological phase transitions mediated by thermal effects. Despite having distinct origins and effects on the band topology, both types of spin-phonon coupling play crucial roles in shaping the topological features of a kagome antiferromagnet.
At a fundamental level, our observations give insights into the optical phonon contribution in a topologically protected frustrated magnetic system. While our analyses are based on a specific model, the potential for experimental realization of such a system enhances the significance of this study.
@@ A.1 Decoupling magnon-phonon modes by canonical transformation
As mentioned, we consider two distinct types of spin-phonon interactions, where the local magnons interact with both the local and non-local optical phonon modes. To incorporate their effects in the effective magnon Hamiltonian, we employ a canonical spin-Peierls transformation [95], where the transformed Hamiltonian reads as,
$\tilde{H}=e^{R^{l(nl)}}He^{-R^{l(nl)}},$ (18)
where $R^{l(nl)}$ is the generator corresponding to the local (non-local) spin-phonon interaction, and can be expressed as,
Figure 10: A single unit cell of the kagome lattice is shown, where each lattice site is labeled with its position coordinate, and the circular spots in grey represent the phonon modes.
@@
Here, $\tilde{\beta}_{i,j}=\tilde{b}_{i}-\tilde{b}_{j}$ and the explicit forms of $R^{l(nl)}$ can be obtained from the HP expansion of $\mathbf{S}_{i}\cdot\mathbf{S}_{j}$. Moreover, Eq. (18) can be cast as a series sum as,
$\tilde{H}=H+[R^{l(nl)},H]+\frac{1}{2!}[R^{l(nl)},[R^{l(nl)},H]]+\cdots.$ (20)
Here, $R^{l(nl)}$ satiesfies the relation $[R^{l(nl)},H_{ph}]=-H_{sp}^{l(nl)}$. Further, it is pertinent to note that in the regime $\hbar\omega\ll J_{1}S$, that is at very low frequencies, the factor $e^{-R^{l(nl)}}$ vanishes. Consequently, the canonical transformation becomes inapplicable in this limit.
Further, this types of spin-phonon couplings give rise to various types of magnon phonon interaction terms, such as, $\tilde{b}_{i}^{(\dagger)}m_{i}^{\dagger}m_{i}$, $\tilde{b}_{i}^{(\dagger)}m_{i}^{\dagger}m_{j}$, $\tilde{b}_{i}^{(\dagger)}m_{i}^{\dagger}m_{j}^{\dagger}$ ($m\in a,b,c$ magnon operators corresponding to the three sublattices), etc. Among these, while the onsite coupling terms provide the dominant contribution, the offsite terms lead to scattering processes whose contributions cannot be captured within a well-defined perturbative expansion using non-interacting formalisms. Therefore, as long as we are working within a non-interacting framework, the off-site contributions, both from the spin-phonon interaction Hamiltonian and the corresponding decoupling operators $R^{l(nl)}$, can be safely neglected. Further, we shall show the final form of the decoupling operators below.
In the presence of phonon operators, different lattice points in a unit cell should be marked with different position indices. Hence, in Fig. (10), we introduce new dummy index $\tilde{r}$ and the NN lattice vectors, $\tilde{\delta}_{i}=\delta_{i}/2$($i\in 1,2,3$). Further, the generators in terms of the magnon and phonon bases can be written as,
@@
where $\alpha^{l(nl)}=(g^{l(nl)}/\hbar\omega)S\mathcal{E}$. Incorporating this into Eq. (20), we obtain the generalized form of the modified Hamiltonian $\tilde{H}$ in terms of the general magnon basis $m$ ($m\in\{a,b,c\}$), which is given by,
$\begin{split}\tilde{H}=S&\left[\sum_{i}\left(4(J_{1}+J_{2})\,{\mathcal{E}}+4D\,{\mathcal{E}}_{1}+\mu_{B}B_{0}\sin{\eta}-\Delta^{l(nl)}\right)m_{i}^{\dagger}m_{i}+\hbar\omega\sum_{i}\tilde{b}_{i}^{\dagger}\tilde{b}_{i}\right.\\ &\left.+\sum_{\langle ij\rangle}\left(J_{1}(P+iQ)-D(P_{1}+iQ_{1})\right)m_{i}^{\dagger}m_{j}e^{X_{1}^{l(nl)}}+\left(J_{1}R-DR_{1}\right)m_{i}^{\dagger}m_{j}^{\dagger}e^{X_{2}^{l(nl)}}\right.\\ &\left.+\sum_{\langle\langle ij\rangle\rangle}\left(J_{2}(P+iQ)\right)m_{i}^{\dagger}m_{j}e^{X_{3}^{l(nl)}}+J_{2}Rm_{i}^{\dagger}m_{j}^{\dagger}e^{X_{4}^{l(nl)}},\right]\end{split}$ (22)

### v2

@@ III.3 Magnon-polaron thermal Hall conductivity
Additionally, in Fig. 7(c), we also illustrate the direct dependence of thermal Hall conductivity ($\kappa_{xy}$) on temperature at three distinct values of $g^{l}$ (a $g^{l}=0$ is included for comparison). In the absence of spin-phonon coupling, the Hall conductivity arises purely from magnons. With no spin-phonon coupling, at low temperatures, $\kappa_{xy}$ acquires a positive contribution and reaches a local maximum, beyond which it decreases and eventually becomes negative. It may be noted that, at a finite temperature, $\kappa_{xy}$ vanishes due to the mutual cancellation of contributions from all the bands. Such behavior of the thermal Hall conductivity is well-established in the literature [103, 20]. Additionally, at very high temperatures, $\kappa_{xy}$ tends to saturate to a finite value.
@@
Figure 8: (a) The Chern number ($C_{n}$) plot corresponding to the lower ($n=1$) and the middle ($n=2$) bands is depicted as a function of $\widetilde{T}$ ($\equiv k_{B}T/J_{1}S$), where within the dotted region, there occurs closure of gaps between the respective bands. (b) The bulk band structure illustrates the gap closing transition. The Berry curvature plots corresponding to the lower band, (c) before and (d) after the gap closing transition, are shown in the $k_{x}$-$k_{y}$ plane.
Figure 9: The behavior of the thermal Hall conductivity ($\kappa_{xy}$) in the presence of non-local spin-phonon interaction is demonstrated. (a) $\kappa_{xy}$ is plotted as a function of external magnetic field for different values of the coupling strength $g^{nl}$ at $\widetilde{T}=0.4$. The transition lines $\bar{s}^{1},~{}\bar{s}^{2}$ denote $g^{nl}/J_{1}=0.4$, while $\bar{s}^{3},~{}\bar{s}^{4}$ refer to $g^{nl}/J_{1}=0.8$, and the $s^{0}$ transition line corresponds to a topological phase transition for all the cases. (b) $\kappa_{xy}$ is presented as a function of $g^{nl}$ at two different temperatures, keeping $B_{0}=0.4B_{s}$, where $\bar{v}^{1}$ and $\bar{v}^{2}$ correspond to $\widetilde{T}=0.4$ and $0.8$, respectively. Additionally, (c) shows the thermal Hall conductivity as a function of temperature for distinct values of $g^{nl}$. Here, $\widetilde{\kappa}_{xy}=\kappa_{xy}/J_{1}S$ and $\widetilde{T}=k_{B}T/J_{1}S$. Throughout this analysis, the other parameters are fixed at $J_{2}=0.03J_{1}$, $D=0.045J_{1}$, and $\hbar\omega=J_{1}S$.
In the presence of local spin-phonon coupling ($g^{l}\neq 0$), the Hall conductivity corresponding to larger values of $g^{l}$ shows a substantial enhancement at low temperatures, whereas at higher temperatures, the growth becomes comparatively less pronounced as seen for $g^{l}=0.8J_{1}$ compared to $g^{l}=0.4J_{1}$ in Fig. 7(c). While in the absence of spin-phonon coupling, the magnon energy bands are independent of temperature, in its presence, the magnon-polaron states are no longer independent of temperature. As temperature increases, we observe positive energy shifts corresponding to the two lower bands (see Appendix C), which affects the $c_{2}(\rho_{n,\mathbf{k}})$ function in such a way that the overall contribution from $c_{2}(\rho_{n,\mathbf{k}})$ function along with temperature responsible for the reduced slope in the behaviour of the thermal Hall conductivity beyond a certain range as it increases. Notably, at higher values of the coupling strength, this behaviour is more prominent. This type of phenomenon reflects the suppression of thermal Hall conductivity due to phonons at high temperatures.
Intriguingly, for $g^{l}=0.4J_{1}$, we find a topological phase transition from $C=(1,-2,1)$ to $C=(-1,0,1)$ occurring at the transition line $T_{c}$. For clarity, the corresponding changes in the Chern numbers of the participating bands are presented in Fig. 8(a) as a function of temperature ($\widetilde{T}$). Moreover, we observe that the effects of temperature are sufficiently weak such that the gap closing transition at the $\mathbf{K}$ point (see Fig. 8(b)) extends over a broad temperature range. This region is highlighted by the dotted area in Fig. 8(a), signifying a weakly semimetalic phase of the bosonic quasiparticles. Owing to the gradual nature of the transition and the fact that it occurs at the $\mathbf{K}$ point, where the contribution from $c_{2}(\rho_{n,\mathbf{k}})$ is relatively small, the transition is not prominently noticeable from the Hall conductivity plot in Fig. 7(c). Furthermore, in Fig. 8(c), for the lower band, we show the corresponding Berry curvature prior to the gap closing transition, where positive values of the Berry curvature are concentrated at the high-symmetry $\mathbf{K}$ and $\mathbf{K^{\prime}}$ points, and they become negative beyond the gap closing transition in Fig. 8(d). This indicates a topological phase transition denoted by $C=1\to C=-1$ corresponding to the lower band.
Similar to the local interaction case, non-local spin-phonon interaction can also lead to some interesting topological phase transitions. The variation of the $\kappa_{xy}$ can be studied as a function of the external magnetic field ($B_{0}$) for two representative values of $g^{nl}$, while keeping the other parameters fixed at $J_{2}=0.03J_{1}$, $D=0.045J_{1}$, $\widetilde{T}=0.4$. For a direct comparison, we also include the results for the bare magnon case ($g^{l(nl)}=0$), as shown in Fig. 9(a). As displayed here, at finite $g^{nl}$, namely, $g^{nl}=0.4J_{1}$, a gradual increase in $B_{0}$ leads to a phase transition from $C=(3,-1,-2)$ to $C=(3,-4,1)$ across the $\bar{s}^{1}$ transition line at $B_{0}=0.12B_{s}$, where the Hall conductivity reaches its peak value. A further increase in $B_{0}$ induces another phase transition at $\bar{s}^{2}$ at $B_{0}=0.25B_{s}$, where the Hall conductivity monotonically decreases. A similar sequence of phase transitions is observed for $g^{nl}=0.8J_{1}$, as shown in Fig. 9(a), where the transition lines $\bar{s}^{3}$ and $\bar{s}^{4}$ are obtained for $B_{0}=0.18B_{s}$ and $B_{0}=0.3B_{s}$, respectively. Moreover, in contrast to the earlier scenario involving local spin-phonon interactions, where the Hall conductivity for different values of $g^{l}$ changes sign (from positive to negative) at distinct value of $B_{0}$, the case corresponding to the non-local interaction exhibits nearly a constant $B_{0}$ at which the Hall conductivity vanishes for $g^{nl}=0.4J_{1}$ and $0.8J_{1}$ (see the clustering of the points on the horizontal dashed line). Consequently, near the $s^{0}$ transition line, another phase transition from $C=(1,-2,1)$ to $C=(-1,0,1)$ occurs for the cases with $g^{nl}=0.4J_{1}$ and $0.8J_{1}$, as shown in Fig. 9(a).
It is evident in Fig. 5(a) that at $\widetilde{T}=0$, with the same values of $D,~{}J_{2}$, and $B_{0}$ mentioned earlier, one phase transition occurs from $C=(3,-4,1)$ to $C=(1,-2,1)$ as a function of $g^{nl}$ near $g^{nl}=0.43J_{1}$. Hence, in Fig. 9(b), we show the corresponding nature of the thermal Hall conductivity as a function of $g^{nl}$ at two different temperatures, where we observe a slight non-monotonic behaviour of the Hall conductivity associated with that phase transition across the transition line $\bar{v}^{1}$ (at $g^{nl}=0.35J_{1}$) and $\bar{v}^{2}$ (at $g^{nl}=0.31J_{1}$) corresponding to the temperatures $\widetilde{T}=0.4$ and $0.5$, respectively. Moreover, it reflects how the transition points shift with varying temperature. Similar to the local case, the Hall conductivity increases with the coupling strength $g^{nl}$ except in the vicinity of the phase transition, as shown in Fig. 9(b).
Additionally, we inspect the behaviour of Hall conductivity as a function of temperature. Here we keep $B_{0}=0.4B_{s}$ same as in Fig. 7(c), and if we see Fig. 9(a), we can observe near $B_{0}=0.4B_{s}$, the behaviour of the Hall conductivity remains nearly same for all the cases corresponding to different $g^{nl}$ which is manifested Fig. 9(c). Moreover, there are a couple of crossings between the Hall conductivity curves corresponding to $g^{nl}=0.4J_{1}$ (in pink) and $0.8J_{1}$ (in blue) as seen in Fig. 9(c). The first crossing at low temperature (near $\widetilde{T}\sim 0.3$) occurs due to a phase transition from $C=(3,-4,1)$ to $C=(1,-2,1)$, corresponding to the case with $g^{nl}=0.4J_{1}$, near the transition line $\overline{T}_{c}$ at $\widetilde{T}=0.29$. A second crossing exists at higher values of $\widetilde{T}$, which however is due to the change in the slope of $\kappa_{xy}$ corresponding to the case with $g^{nl}=0.8J_{1}$, which was earlier denoted as the suppression of the thermal Hall conductivity at high temperature induced by phonons.
Figure 10: (a) The Chern number ($C_{n}$) plot corresponding to the lower ($n=1$) and the middle ($n=2$) bands is displayed as a function of $\widetilde{T}$ ($\equiv k_{B}T/J_{1}S$), where within the dotted region, there occurs closure of gaps between the respective bands, which is further illustrated in (b) via the gap closing transition. The Berry curvature plots corresponding to the lower band, (c) before and (d) after the gap closing transition, are shown in the $k_{x}$-$k_{y}$ plane.
@@
To clarify the above mentioned phase transition for $g^{nl}=0.4J_{1}$, we further present the corresponding Chern number plot as a function of temperature in Fig. 10 (a), where the dotted region corresponds to the spectral gap closing regime. Further, in Fig. 10(b), we observe a gap closing transition near the $\mathbf{\Gamma}$ point, where the contribution from $c_{2}(\rho_{n,\mathbf{k}})$ is significant at low temperatures. As a result, the transition appears more prominently in the Hall conductivity plot compared to the case with local spin-phonon interaction. Additionally, in Fig. 10(c), at $\widetilde{T}=0.24$, prior to the gap closing transition, we show the Berry curvature of the lower band. Positive values of the Berry curvature are observed near both the high-symmetry points ($\mathbf{\Gamma}$ and $\mathbf{K}$), corresponding to a Chern number of $C=3$. In contrast, in Fig. 10(d), after closing the spectral gap, the Berry curvature changes sign (negative) at the gap closing point $\mathbf{\Gamma}$, while remaining positive at the $\mathbf{K}$ point, indicating $C=1$.
@@ IV CONCLUSION
We investigate the effects of couplings between spins and optical phonons in a topologically protected magnetic system, namely, in a frustrated Kagome antiferromagnet that attains a canted configuration under the application of an external magnetic field. Specifically, we consider two different types of spin-phonon couplings, one in which the interaction is purely local, while the other accounts for non-local interactions, to encounter and compare different kinds of topological scenarios. To deal with the LO phonons, we perform detailed and systematic calculations based on a canonical spin-Peierls transformation formulated within the magnon framework and obtain a hybridized magnon-polaron state. Interestingly, we observe several topological phase transitions solely induced by the spin-phonon couplings, which are explored via the magnon-polaron band structures, their Chern numbers, and the corresponding phase plots. Furthermore, the thermal Hall conductivities relevant to these scenarios are obtained, thereby providing robust support for these topological transitions.
To enumerate the highlights of our results, we observe multiple topological phases accompanied by several bulk gap closing transitions as a function of the DMI and the spin-phonon coupling strength, albeit demonstrating distinct signatures for the local and non-local spin-phonon couplings. To establish the bulk-boundary correspondence, we show the variations of the associated edge modes on a ribbon geometry, where these distinct phases are validated via computing the winding number, which dictates the number of edge crossings and the chirality of the magnon-polaron edge currents. We emphasize that for an identical range of the coupling strengths, the local coupling mechanism renders a larger number of distinct topological phases, compared to the non-local one. Furthermore, in this study, we observe that due to the inclusion of the phonon contributions at finite temperatures, the spectral properties of the energy bands depend on temperature, resulting in topological phase transitions mediated via thermal effects. Moreover, at finite temperatures, the behaviour of the thermal Hall conductivity as a function of the external magnetic field, spin-phonon coupling strength, and temperature, effectively captures all the topological phases and phase transitions occurring therein. The contrasting effects resulting from the two types of couplings further corroborate that local coupling exhibits a comparatively stronger effect on enhancing the thermal Hall conductivity than that arising out of the non-local coupling (particularly when the Hall conductivity nearly vanishes for the bare magnon case). Despite having distinct origins and effects on the topology of the magnon-polaron bands, both types of spin-phonon coupling play crucial roles in shaping the topological features of a frustrated kagome antiferromagnet.
At a fundamental level, our observations give insights into the optical phonon contribution in a topologically protected frustrated magnetic system. While our analyses are based on a specific model, the potential for experimental realization of such a system strengthens the significance of this study.
Acknowledgments
S.D. acknowledges financial support from the Ministry of Education (MOE), Govt. of India. K.B. acknowledges financial support from the Anusandhan National Research Foundation (ANRF), Govt. of India, through the National Post Doctoral Fellowship (NPDF) (File No. PDF/2023/000161).
@@ A.1 Decoupling magnon-phonon modes by canonical transformation
As mentioned, we consider two distinct types of spin-phonon interactions, where the local magnons interact with both the local and non-local optical phonon modes. To incorporate the corresponding effects in the effective magnon Hamiltonian, we employ a canonical spin-Peierls transformation [95], via the operator $U^{l(nl)}=e^{R^{l(nl)}}$, where the transformed Hamiltonian reads as
$\widetilde{H}=e^{R^{l(nl)}}He^{-R^{l(nl)}},$ (18)
and the corresponding hybrid wavefunction of the system can be represented as a tensor product of magnonic and phononic wavefunctions, which looks like
$|\phi\rangle_{\text{hyb}}=U^{l(nl)}|\phi\rangle_{\text{ph}}\otimes|\phi\rangle_{\text{mag}}.$ (19)
The unitary generator of the transformation, namely $R^{l(nl)}$ corresponding to the local (non-local) spin-phonon interaction, can be expressed as
@@
Here, $\tilde{\beta}_{i,j}=\tilde{b}_{i}-\tilde{b}_{j}$ and the explicit forms of $R^{l(nl)}$ can be obtained from the HP expansion of $\mathbf{S}_{i}\cdot\mathbf{S}_{j}$. Moreover, Eq. (18) can be cast as a series sum and is given by
$\widetilde{H}=H+[R^{l(nl)},H]+\frac{1}{2!}[R^{l(nl)},[R^{l(nl)},H]]+\cdots.$ (21)
Figure 11: A schematic representation of a generic single unit cell of a KAFM lattice is shown, where each lattice site is labeled with its position coordinate, namely $\tilde{r}\mp\tilde{\delta}_{3}$, $\tilde{r}$, and $\tilde{r}\mp\tilde{\delta}_{2}$ for A, B, and C sublattices, respectively with $\tilde{\delta}_{i}$ ($i\in 1,2,3$) denoting three NN lattice vectors. The circular spots in grey represent the phonon modes corresponding to each site.
Here, $R^{l(nl)}$ satisfies the relation $[R^{l(nl)},H_{\text{ph}}]=-H_{\text{sp}}^{l(nl)}$. Further, it is pertinent to note that in the regime $\hbar\omega\ll J_{1}S$, that is at very low frequencies, the factor $e^{-R^{l(nl)}}$ vanishes. Consequently, the canonical transformation becomes inapplicable in this limit.
Further, this types of spin-phonon couplings give rise to various types of magnon phonon interaction terms, such as, $\tilde{b}_{i}^{(\dagger)}m_{i}^{\dagger}m_{i}$, $\tilde{b}_{i}^{(\dagger)}m_{i}^{\dagger}m_{j}$, $\tilde{b}_{i}^{(\dagger)}m_{i}^{\dagger}m_{j}^{\dagger}$ ($m\in a,b,c$ corresponding to the three sublattices), etc. Among these, while the onsite coupling terms provide the dominant contribution, the offsite terms lead to scattering processes whose contributions cannot be captured within a well-defined perturbative expansion using non-interacting formalisms. Therefore, as long as we are working within a non-interacting framework, the off-site contributions, both from the spin-phonon interaction Hamiltonian and the corresponding decoupling operators $R^{l(nl)}$, can be safely neglected. Further, we shall show the final form of the decoupling operators below.
In the presence of phonon operators, different lattice points in a unit cell should be marked with different position indices. Hence, in Fig. (11), we introduce new dummy index $\tilde{r}$ and the NN lattice vectors, $\tilde{\delta}_{i}=\frac{\delta_{i}}{2}$ ($i\in 1,2,3$). Further, the generators in terms of the magnon and phonon bases can be written as
@@
where $\alpha^{l(nl)}=(g^{l(nl)}/\hbar\omega)S\mathcal{E}$. Incorporating this into Eq. (21), we obtain the generalized form of the modified Hamiltonian $\widetilde{H}$ in terms of the general magnon basis $m$ ($m\in\{a,b,c\}$), which is given by
$\begin{split}\widetilde{H}=S&\left[\sum_{i}\left(4(J_{1}+J_{2})\,{\mathcal{E}}+4D\,{\mathcal{E}}_{1}+\mu_{B}B_{0}\sin{\eta}-\Delta^{l(nl)}\right)m_{i}^{\dagger}m_{i}+\hbar\omega\sum_{i}\tilde{b}_{i}^{\dagger}\tilde{b}_{i}\right.\\ &\left.+\sum_{\langle ij\rangle}\left(J_{1}(P+iQ)-D(P_{1}+iQ_{1})\right)m_{i}^{\dagger}m_{j}e^{X_{1}^{l(nl)}}+\left(J_{1}R-DR_{1}\right)m_{i}^{\dagger}m_{j}^{\dagger}e^{X_{2}^{l(nl)}}\right.\\ &\left.+\sum_{\langle\langle ij\rangle\rangle}\left(J_{2}(P+iQ)\right)m_{i}^{\dagger}m_{j}e^{X_{3}^{l(nl)}}+J_{2}Rm_{i}^{\dagger}m_{j}^{\dagger}e^{X_{4}^{l(nl)}},\right]\end{split}$ (23)